OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x2 – 10x + 8? solving quadratics by completing the square

OpenStudy (anonymous):

plug x=0 for y int plug y=0 for x int

OpenStudy (anonymous):

i dont get it?

jimthompson5910 (jim_thompson5910):

x-intercepts y = 2x^2 – 10x + 8 0 = 2x^2 – 10x + 8 0 = 2(x^2 – 5x + 4) 0 = 2(x-1)(x-4) x-1=0 or x-4=0 x=1 or x=4 x-intercepts are (1,0) and (4,0) ---------------------------- y-intercept y = 2x^2 – 10x + 8 y = 2(0)^2 – 10(0) + 8 y = 8 y-intercept is (0,8)